# Scalable and Accurate Graph Reasoning with LLM-based Multi-Agents

### v2

@@ Abstract
Recent research has explored the use of Large Language Models (LLMs) for tackling complex graph reasoning tasks. However, due to the intricacies of graph structures and the inherent limitations of LLMs in handling long text, current approaches often fail to deliver satisfactory accuracy, even on small-scale graphs and simple tasks. To address these challenges, we introduce GraphAgent-Reasoner, a fine-tuning-free framework that utilizes a multi-agent collaboration strategy for explicit and precise graph reasoning. Inspired by distributed graph computation theory, our framework decomposes graph problems into smaller, node-centric tasks that are distributed among multiple agents. The agents collaborate to solve the overall problem, significantly reducing the amount of information and complexity handled by a single LLM, thus enhancing the accuracy of graph reasoning. By simply increasing the number of agents, GraphAgent-Reasoner can efficiently scale to accommodate larger graphs with over 1,000 nodes. Evaluated on the GraphInstruct dataset, our framework demonstrates near-perfect accuracy on polynomial-time graph reasoning tasks, significantly outperforming the best available models, both closed-source and fine-tuned open-source variants. Our framework also demonstrates the capability to handle real-world graph reasoning applications such as webpage importance analysis.
@@ 1 Introduction
Despite significant efforts by researchers to enable LLMs to memorize, comprehend, and perform basic reasoning on graph structures, several issues still persist: 1) The scale of graphs that can be handled is limited. Describing graph structures in natural language inevitably leads to excessively long inputs. Due to context length limitations and the shortcomings of LLMs in handling lengthy text (Liu2023LostIT), previous works (chai2023graphllm; fatemi2023talk; perozzi2024let) could only handle graphs of very limited size (e.g. fewer than 20 nodes and 100 edges). 2) The performance on graph reasoning tasks is relatively poor. Unlike text, which can tolerate some degree of semantic deviation, reasoning and computation on graphs must be highly precise. However, current works demonstrate poor accuracy (average 2060%) in various graph reasoning tasks like connectivity and shortest path. 3) Lacking explicit reasoning paths. Taking the shortest path as an example, the responses of existing models resemble a heuristic search approach to finding the shortest path on a graph, rather than strictly executing an algorithm. This makes it difficult to determine whether LLMs are genuinely deriving the answer through correct reasoning or merely making educated guesses. Although GraphWiz (chen2024graphwiz) attempts to generate explicit reasoning paths through fine-tuning, it often fails due to the presence of incomplete or wrong reasoning paths in its training data. Furthermore, GraphWiz exhibits overfitting, where it tends to treat new or unrelated questions as one of the fine-tuned problems, which will be detailed in Section 5.3.
@@
Motivation. The ultimate goal of graph reasoning is to enable LLMs to leverage graph-related knowledge or algorithms to solve real-world graph problems. However, with the development of information science and hardware storage, the scale of graphs and information per node become too large for a single LLM to handle. To address this, a natural idea is to use distributed approaches, where a large graph is stored across multiple LLMs separately and compute collaboratively. Therefore, just as graph algorithms have generally evolved from non-distributed to distributed forms (meng2024survey)), we hope that LLMs can also learn the concept of distributed processing, thereby harnessing the power of swarm intelligence to solve graph problems in real-world scenarios.
@@
Our Contribution. To address the above limitations, in this paper, we propose the GraphAgent-Reasoner(GAR) framework, which leverages the power of swarm intelligence to solve graph reasoning problems, as shown in Figure 1. We follow a node-centric approach, assigning an agent to each node, allowing it to focus on processing its own information and communicate with neighbors. Thus, we can easily scale up the size of graphs that can be processed by simply increasing the number of agents. At the same time, under the direction of a Master LLM, graph problems are decomposed into smaller, node-centric tasks, which are assigned to agents for collaborative resolution. This approach significantly reduces the scale and complexity of information each agent needs to process, thereby greatly improving the overall accuracy. Furthermore, since agents must clearly transmit the processed information to neighboring agents, the reasoning process becomes transparent, demonstrating the framework solves graph reasoning problems through clear and correct reasoning, rather than lucky guessing. In summary, our contributions are as follows:
@@
We propose GraphAgent-Reasoner, the first LLM-based multi-agents framework for graph reasoning, which requires no fine-tuning and can utilize any LLM as the underlying reasoning model. Our framework achieves near-perfect accuracy on various polynomial-time tasks, significantly surpassing the performance of existing methods.
@@ 2 Preliminaries and Related Works
Preliminaries. In general scenarios, when discussing LLMs solving graph reasoning problems, the input is a (,) pair. is a graph represented as , where is the node set and , the edge set. For each node , a sequential text node feature is associated; similarly, for each edge , a sequential text edge feature is assigned. The graph is described in natural language, typically using edge or adjacency list representation. is a task-specific instruction or problem description. LLMs will process the (,) pair and return an answer string .
Large Language Models for Graph Reasoning. To further enhance the reasoning capabilities of LLMs, many works have attempted to improve the performance of LLMs in graph reasoning. wang2024can first introduces the NLGraph Benchmark to evaluate the performance of LLMs on various graph reasoning tasks. fatemi2023talk explores the impact of different graph encoding methods and graph structure types on the performance of LLMs in graph reasoning tasks. Additionally, it introduces another benchmark called GraphQA. Considering the lengthy nature of describing graph structures in text, chai2023graphllm and perozzi2024let respectively use Transformers and GNNs to encode graph structures and attempt to align them with LLMs. Inspired by how humans understand structural information through the visual modality, Wei2024GITAGT generates corresponding visual images based on graph structures and provides them to visual LLMs for graph reasoning. chen2024graphwiz conducted Supervised Fine-Tuning and Directly Prefered Optimization on LLMs, enhancing the performance of LLMs and encouraging them to output explicit reasoning paths.
Large Language Model based Multi-Agents. Recent advancements in LLMs have spurred interest in their application within multi-agent systems. LLM-based multi-agent frameworks leverage the natural language understanding and reasoning capabilities of LLMs to enable agents to collaborate, communicate, and solve complex tasks in a distributed manner. Existing multi-agents works for problem solving primarily focuses on applications such as Software Development (dong2023self; hong2024metagpt; qian2024chatdev), Embodied Agents (zhang2024build; zhao2024roco; chen2024scalable) and Science Debate (xiong2023examing; chan2024chateval). However, using LLM-based multi-agents to handle graph data has been less explored, especially in the areas of graph reasoning and graph computation tasks. This may be due to the hallucination issue inherent in LLMs (huang2023hallucination), where their responses are factually incorrect. This problem becomes more complex in a multi-agent setting, as the hallucinations of a single agent may propagate to other nodes by communication (guo2024multiagent). This requires the performance of individual agents be sufficiently stable to ensure the correct operation of the entire multi-agent system.
3 Limitations of Single LLM in Graph Reasoning
@@
The graph structure is too complex to memorize and understand for a single LLM. Using adjacency or edge lists to describe graph structures in natural language is the most intuitive and direct method, facilitating the processing of graph data by LLMs through text. However, this approach inevitably leads to a lengthy context, as the number of edges can grow quadratically with the number of nodes. As the graph scales up and becomes denser, the graph structure becomes highly complex, requiring a large amount of tokens to describe the edge relationships. When the text becomes too lengthy, it becomes difficult for LLMs to properly allocate attention, and they may even struggle with simple tasks such as key-value pair matching Liu2023LostIT. This presents significant challenges for LLMs in identifying key information for graph reasoning tasks from the lengthy context. Figure 2 shows the performance of a single LLM in memorizing one-hop neighbor nodes. We observe that as the number of nodes in the graph increases, various LLMs exhibit a significant decline in accuracy. If a single LLM cannot even correctly recall basic graph structural information like node neighbors, it becomes difficult to proceed with more complex graph reasoning or computation.
@@
A single LLM struggles to solve reasoning problems in real-world scenarios. Researchers train LLMs on graph reasoning tasks to empower them to utilize learned graph-related knowledge or algorithms to tackle real-world graph problems. However, in practical scenarios, the amount of information associated with each node can be enormous. Take citation networks as an example: a single node represents a paper, and its node information includes the title, abstract, and references, which could amount to several thousand tokens. In addition to the complexity of graph structures, the need to handle a large amount of node information further exacerbates the burden on a single LLM and highlights its shortcomings in processing long contexts. Moreover, using a single LLM to handle the entire network is inefficient, as it cannot coherently process the entire network’s problems. Typically, it is necessary to manually compress or summarize the information for each node and then feed local subgraphs to the LLM for processing (guo2023gpt4graph; chen2024exploring).
@@
4 GraphAgent-Reasoner
To solve the limitations above, we propose a novel framework based on multi-agent collaboration called GraphAgent-Reasoner as shown in Figure 3, aiming to solve graph reasoning problems explicitly and correctly. The interface of the framework is a Master LLM, which is responsible for processing the textual input of graph problems, constructing the agent network, directing them to collaboratively solve the problem, and finally aggregating the states of all agents to derive the solution. Its implementation is based on the React Agent proposed by yao2023react, which is capable of reasoning based on the environment and executing corresponding actions, as detailed later. The pipeline of GAR consists of four steps: Graph Construction, Algorithm Establishing, Distributed Execution and Master Summarization.
Graph Construction. Given an input pair (, ), the Master LLM first extracts the node and edge information from the textual description of graph . It then constructs an agent for each node and initializes the node’s state and neighbor information, forming an interconnected network of agents. Each agent independently maintains its state and neighbor data, communicates with adjacent agents based on instructions from the Master LLM, and updates its state in each round.
Algorithm Establishing. To accommodate diverse graph tasks and fully exploit the knowledge embedded in LLMs during pre-training, we propose a unified solution approach framed within a distributed paradigm as shown in Algorithm 1. This approach requires the Master LLM to specify six core components for each problem: State, Message, Initialization, Send, Update, and Termination.
State: The local information maintained by each node, representing its current state. This can include attributes like node features, labels, or any other task-specific data. The states evolve as nodes receive messages and update their information.
Message: The data transmitted between nodes during the communication phase. Messages typically contain information that neighboring nodes need to perform updates, such as feature values, distances, or other task-relevant information.
@@ 4 GraphAgent-Reasoner
Update: Upon receiving messages from its neighbors, each node updates its state by aggregating the incoming messages and combining them with its current state. This iterative process enables nodes to refine their information over time.
@@
Since LLMs lack prior knowledge of this distributed paradigm, to facilitate the Master LLM’s understanding and application of the framework, we develop a distributed algorithm library that adheres to this distributed paradigm, from which the Master LLM can query relevant algorithm templates to generate distributed solutions within this paradigm. Specifically, we selected classic distributed graph algorithms and documented their implementations under this distributed paradigm. Some examples are presented in Appendix A.1. Drawing on prior work (zheng2024executing; jojic2024gpt), we endeavor to write detailed reasoning steps of each part in the algorithm to encourage the agent to think step by step as much as possible, which plays an important role in enhancing the success rate of individual agents.
When receiving a problem input, the Master LLM first retrieves the algorithms most relevant to the problem description from the distributed algorithm library. If there are algorithms suitable for handling the problem, the Master LLM will adjust the algorithm according to the problem description, such as changing the initialization and termination conditions (e.g., the source node in the shortest path problem). If there are no appropriate algorithms, the Master LLM will design a distributed algorithm following the distributed paradigm based on the examples of the retrieved algorithms. For some generated examples, see Appendix A.2.
Distributed Execution. After the distributed algorithm is designed, the Master LLM will relay the approach to each agent node for execution according to the process outlined in Algorithm 1. Each agent will first initialize its state based on node information and algorithm rules and then send an initial message to neighboring agents. Subsequently, each agent will iteratively execute the operations of receiving messages, updating its state, and sending messages according to the algorithm rules, synchronizing progress after each communication round. Communication will continue until the maximum number of iterations is reached or the termination condition is met.
Master Summarization. Finally, the final state of all agent nodes will be aggregated to the Master LLM, which will summarize the results conclude based on the problem and return the final answer in natural language form.
@@ 5 Experiments
R1: GAR achieves near-perfect accuracy on polynomial-time graph reasoning problems, significantly surpassing existing closed-source models and open-source models fine-tuned on extensive data.
@@
Datasets. We conduct our experiments on the graph reasoning tasks proposed in GraphInstruct (chen2024graphwiz). This dataset contains nine graph reasoning problems with different time complexity, ranging from linear and polynomial complexity to NP-complete.
Due to the complexity of NP-complete problems, there are currently no mature exact distributed algorithms available for their solution. Consequently, the Master LLM is unable to design correct and effective distributed algorithms based on the knowledge acquired during pre-training. Therefore, in our experiments, we only consider linear and polynomial-time problems. Detailed information of the dataset and partial test results for NP-complete problems will be presented in Appendix B.
Setting. The underlying reasoning LLM of Agent Node used in our framework is ChatGPT-4o-mini-2024-07-18, and the base model of Master LLM is ChatGPT-4-turbo(achiam2023gpt). The temperature is consistently set to 0. Our framwork is built upon AgentScope (agentscope), an innovative platform to easily build reliable, high-performance multi-agent applications.
5.1 Experiment 1: Performance on GraphInstruct
In this experiment, we evaluate the performance of GAR on polynomial-time tasks of the GraphInstruct dataset. The results are shown in Table 1. We see GAR exhibits near-perfect results on these tasks, significantly outperforming other models. Especially on shortest and triangle tasks with high time complexity, GAR substantially improves the performance of LLMs. Problems that a single LLM struggles to solve have been effectively resolved through collaboration by agents after being decomposed into smaller, node-centric tasks.
@@ 5.1 Experiment 1: Performance on GraphInstruct
5.2 Experiment 2: Performance on Large-Scale Graphs

### v3

@@ Abstract
Recent research has explored the use of Large Language Models (LLMs) for tackling complex graph reasoning tasks. However, due to the intricacies of graph structures and the inherent limitations of LLMs in handling long text, current approaches often fail to deliver satisfactory accuracy, even on small-scale graphs and simple tasks. To address these challenges, we introduce GraphAgent-Reasoner, a fine-tuning-free framework that utilizes a multi-agent collaboration strategy for explicit and precise graph reasoning. Inspired by distributed graph computation theory, our framework decomposes graph problems into smaller, node-centric tasks that are distributed among multiple agents. The agents collaborate to solve the overall problem, significantly reducing the amount of information and complexity handled by a single LLM, thus enhancing the accuracy of graph reasoning. By simply increasing the number of agents, GraphAgent-Reasoner can efficiently scale to accommodate larger graphs with over 1,000 nodes. Evaluated on the GraphInstruct dataset, our framework demonstrates impressive accuracy on polynomial-time graph reasoning tasks, significantly outperforming the best available models, both closed-source and fine-tuned open-source variants. Our framework also demonstrates the capability to handle real-world graph reasoning applications such as webpage importance analysis.
@@ 1 Introduction
Despite significant efforts by researchers to enable LLMs to memorize, comprehend, and perform basic reasoning on graph structures, several issues still persist: 1) The scale of graphs that can be handled is limited. Describing graph structures in natural language inevitably leads to excessively long inputs. Due to context length limitations and the shortcomings of LLMs in handling lengthy text (Liu2023LostIT), previous works (chai2023graphllm; fatemi2023talk; perozzi2024let) could only handle graphs of very limited size (e.g. fewer than 20 nodes and 100 edges). 2) The performance on graph reasoning tasks is relatively poor. Unlike text, which can tolerate some degree of semantic deviation, reasoning and computation on graphs must be highly precise. However, current works demonstrate poor accuracy (average 2060%) in various graph reasoning tasks like connectivity and shortest path. 3) Lacking explicit reasoning paths. Taking the shortest path as an example, the responses of existing models resemble a heuristic search approach to finding the shortest path on a graph, rather than strictly executing an algorithm. This makes it difficult to determine whether LLMs are genuinely deriving the answer through correct reasoning or merely making educated guesses. Although GraphWiz (chen2024graphwiz) attempts to generate explicit reasoning paths through fine-tuning, it often fails due to the presence of incomplete or wrong reasoning paths in its training data. Furthermore, GraphWiz exhibits overfitting, which will be detailed in Section 5.3.
Motivation.
@@
The ultimate goal of graph reasoning is to enable LLMs to leverage graph-related knowledge or algorithms to solve real-world graph problems. However, with the development of information science and hardware storage, both the scale of graphs and the amount of information per node have become too large for a single LLM to handle. To address this, a natural idea is to use distributed approaches, where a large graph is stored across multiple LLMs separately and compute collaboratively. Therefore, just as graph algorithms have generally evolved from non-distributed to distributed forms (meng2024survey)), we hope that LLMs can also learn the concept of distributed processing, thereby harnessing the power of swarm intelligence to solve graph problems in real-world scenarios.
Our Contribution.
@@
To address the above limitations, in this paper, we propose the GraphAgent-Reasoner(GAR) framework, which leverages the power of swarm intelligence to solve graph reasoning problems, as shown in Figure 1. We follow a node-centric approach, assigning an agent to each node, allowing it to focus on processing its own information and communicate with neighbors. Thus, we can easily scale up the size of graphs that can be processed by simply increasing the number of agents. At the same time, under the direction of a Master LLM, graph problems are decomposed into smaller, node-centric tasks, which are assigned to agents for collaborative resolution. This approach significantly reduces the scale and complexity of information each agent needs to process, thereby greatly improving the overall accuracy. Furthermore, since agents must clearly transmit the processed information to neighboring agents, the reasoning process becomes transparent, demonstrating the framework solves graph reasoning problems through clear and correct reasoning, rather than lucky guessing. In summary, our contributions are as follows:
@@
We propose GraphAgent-Reasoner, a LLM-based multi-agents framework for graph reasoning, which requires no fine-tuning and can utilize any LLM as the backbone model. Our framework achieves impressive accuracy on various polynomial-time tasks, significantly surpassing the performance of existing methods.
@@ 2 Preliminaries and Related Works
Preliminaries.
In general scenarios, when discussing LLMs solving graph reasoning problems, the input is a (,) pair. is a graph represented as , where is the node set and , the edge set. For each node , a sequential text node feature is associated; similarly, for each edge , a sequential text edge feature is assigned. The graph is described in natural language, typically using edge or adjacency list representation. is a task-specific instruction or problem description. LLMs will process the (,) pair and return an answer string .
Large Language Models for Graph Reasoning.
To further enhance the reasoning capabilities of LLMs, many works have attempted to improve the performance of LLMs in graph reasoning. wang2024can first introduces the NLGraph Benchmark to evaluate the performance of LLMs on various graph reasoning tasks. fatemi2023talk explores the impact of different graph encoding methods and graph structure types on the performance of LLMs in graph reasoning tasks. Additionally, it introduces another benchmark called GraphQA. Considering the lengthy nature of describing graph structures in text, chai2023graphllm and perozzi2024let respectively use Transformers and GNNs to encode graph structures and attempt to align them with LLMs. Inspired by how humans understand structural information through the visual modality, Wei2024GITAGT generates corresponding visual images based on graph structures and provides them to visual LLMs for graph reasoning. chen2024graphwiz conducted Supervised Fine-Tuning and Directly Prefered Optimization on LLMs, enhancing the performance of LLMs and encouraging them to output explicit reasoning paths.
Large Language Model based Multi-Agents.
Recent advancements in LLMs have spurred interest in their application within multi-agent systems. LLM-based multi-agent frameworks leverage the natural language understanding and reasoning capabilities of LLMs to enable agents to collaborate, communicate, and solve complex tasks in a distributed manner. Existing multi-agents works for problem solving primarily focuses on applications such as Software Development (dong2023self; hong2024metagpt; qian2024chatdev), Embodied Agents (zhang2024build; zhao2024roco; chen2024scalable) and Science Debate (xiong2023examing; chan2024chateval). However, using LLM-based multi-agents to handle graph data has been less explored, especially in the areas of graph reasoning and graph computation tasks. This may be due to the hallucination issue inherent in LLMs (huang2023hallucination), where their responses are factually incorrect. This problem becomes more complex in a multi-agent setting, as the hallucinations of a single agent may propagate to other nodes by communication (guo2024multiagent). This requires the performance of individual agents be sufficiently stable to ensure the correct operation of the entire multi-agent system.
3 Limitations of a Single LLM
@@
The graph structure is too complex to memorize and understand for a single LLM.
Using adjacency or edge lists to describe graph structures in natural language is the most intuitive and direct method, facilitating the processing of graph data by LLMs through text. However, this approach inevitably leads to a lengthy context, as the number of edges can grow quadratically with the number of nodes. As the graph scales up and becomes denser, the graph structure becomes highly complex, requiring a large amount of tokens to describe the edge relationships. When the text becomes too lengthy, it becomes difficult for LLMs to properly allocate attention, and they may even struggle with simple tasks such as key-value pair matching (Liu2023LostIT). This presents significant challenges for LLMs in identifying key information for graph reasoning tasks from the lengthy context. Figure 2 shows the performance of a single LLM in memorizing one-hop neighbor nodes. We observe that as the number of nodes in the graph increases, various LLMs exhibit a significant decline in accuracy. If a single LLM cannot even correctly recall basic graph structural information like node neighbors, it becomes difficult to proceed with more complex graph reasoning or computation.
@@
A single LLM struggles to solve reasoning problems in real-world scenarios.
Researchers fine-tune LLMs on graph reasoning tasks to empower themto enhance their ability to leverage learned graph-related knowledge or algorithms in solving real-world graph problems. However, in practical applications, the volume of information associated with each node can be substantial. Take citation networks as an example: a single node represents a paper, and its node information includes the title, abstract, and references, which could amount to several thousand tokens. In addition to the complexity of graph structures, the need to handle a large amount of node information further exacerbates the burden on a single LLM and highlights its shortcomings in processing long contexts. Moreover, using a single LLM to handle the entire network is inefficient, as it cannot coherently process the entire network’s problems. Typically, it is necessary to manually compress or summarize the information for each node and then feed local subgraphs to the LLM for processing (guo2023gpt4graph; chen2024exploring).
@@
4 Our Approach
To solve the limitations above, we propose a novel framework based on multi-agent collaboration called GraphAgent-Reasoner as shown in Figure 3, aiming to solve graph reasoning problems explicitly and correctly. The interface of the framework is a powerful LLM called Master, which is responsible for processing the textual input of graph problems, constructing the agent network, directing them to collaboratively solve the problem, and finally aggregating the states of all agents to derive the solution. Its implementation is based on the React Agent proposed by yao2023react, which is capable of reasoning based on the environment and executing corresponding actions, as detailed later. The pipeline of GAR consists of four steps: Graph Construction, Algorithm Establishing, Distributed Execution and Master Summarization.
Graph Construction.
Given an input pair (, ), the Master first extracts the node and edge information from the textual description of graph . It then constructs an agent for each node and initializes the node’s state and neighbor information, forming an interconnected network of agents. Each agent independently maintains its state and neighbor data, communicates with adjacent agents based on instructions from the Master, and updates its state in each round.
Algorithm Establishing.
To accommodate diverse graph tasks and fully exploit the knowledge embedded in LLMs during pre-training, we propose a unified solution approach framed within a distributed paradigm as shown in Algorithm 1. This approach requires the Master LLM to specify six core components for each problem: State, Message, Initialization, Send, Update, and Termination.
State: The local information maintained by each node, representing its current state. This can include attributes like node features, labels, long texts or any other task-specific data. The states evolve as nodes receive messages and update their information.
Message: The information exchanged between neighboring nodes during communication rounds. Messages typically contain information that neighboring nodes need to perform updates, such as feature values, distances, or other task-relevant information.
@@ 4 Our Approach
Update: Upon receiving messages from its neighbors, each node updates its state by aggregating the incoming messages and combining them with its current state. This iterative process enables nodes to update their local information over time.
@@
Input: Agent Nodes , each agent maintains a state , the maximum iterations given by the Master LLM
Output: Final state for each agent
Since LLMs lack prior knowledge of this distributed paradigm, to facilitate the Master LLM’s understanding and application of the framework, we develop a distributed algorithm library that adheres to this distributed paradigm, from which the Master LLM can query relevant algorithm templates to generate distributed solutions within this paradigm. Specifically, we selected classic distributed graph algorithms and documented their implementations under this distributed paradigm. Some examples are presented in Appendix 8.3. Drawing on prior work (zheng2024executing; jojic2024gpt), we endeavor to write detailed reasoning steps of each part in the algorithm to encourage the agent to think step by step as much as possible, which plays an important role in enhancing the success rate of individual agents.
When receiving a problem input, the Master LLM first retrieves the algorithms most relevant to the problem description from the distributed algorithm library. If there are algorithms suitable for handling the problem, the Master LLM will adjust the algorithm according to the problem description, such as changing the initialization and termination conditions (e.g., the source node in the shortest path problem). If there are no appropriate algorithms, the Master LLM will design a distributed algorithm following the distributed paradigm based on the examples of the retrieved algorithms. For some generated examples, see Appendix 8.3.
Distributed Execution.
After the distributed algorithm is designed, the Master LLM will relay the approach to each agent node for execution according to the process outlined in Algorithm 1. Each agent will first initialize its state based on node information and algorithm rules and then send an initial message to neighboring agents. Subsequently, each agent will iteratively execute the operations of receiving messages, updating its state, and sending messages according to the algorithm rules, synchronizing progress after each communication round. Communication will continue until the maximum number of iterations is reached or the termination condition is met.
Master Summarization.
Finally, the final state of all agent nodes will be aggregated to the Master LLM, which will summarize the results conclude based on the problem and return the final answer in natural language form.
@@ 5 Experiments
R1: GAR achieves impressive accuracy on polynomial-time graph reasoning problems, significantly surpassing existing closed-source models and open-source models fine-tuned on extensive data.
@@
Datasets.
We conduct our experiments on the graph reasoning tasks proposed in GraphInstruct (chen2024graphwiz). This dataset contains nine graph reasoning problems with different time complexity, ranging from linear and polynomial complexity to NP-complete.
Due to the complexity of NP-complete problems, there are currently no mature exact distributed algorithms available for their solution. Consequently, the Master LLM is unable to design correct and effective distributed algorithms based on the knowledge acquired during pre-training. Therefore, in our experiments, we only consider linear and polynomial-time problems. Detailed information of the dataset and partial test results for NP-complete problems will be presented in Appendix 8.1.
Setting.
The underlying reasoning LLM of Agent Node used in our framework is ChatGPT-4o-mini-2024-07-18, and the base model of Master LLM is ChatGPT-4-turbo(achiam2023gpt). The temperature is consistently set to 0. Our framwork is built upon AgentScope ((agentscope)), an innovative platform to easily build reliable, high-performance multi-agent applications.
5.1 Performance on GraphInstruct
In this experiment, we evaluate the performance of GAR on polynomial-time tasks of the GraphInstruct dataset. The results are shown in Table 1. We see GAR exhibits impressive results on these tasks, significantly outperforming other models. Especially on shortest and triangle tasks with high time complexity, GAR substantially improves the performance of LLMs. Problems that a single LLM struggles to solve have been effectively resolved through collaboration by agents after being decomposed into smaller, node-centric tasks.
5.2 Performance on Large-Scale Graphs